Write the percent of the number.
10. 50% of 22
Any help with this equation would be Greatly appreciated. Answer by keerthu(5) 
 10.50 % of 22 = (10.50/100)*22
= 0.1050 * 22
= 2.31
Just now i came to know that 10 is the question number and not a part of question...
Hence, the answer goes like this...
50% of 22 = (50/100)*22
= 0.50 * 22
= 11
Devnull, hope your problem is solved.